OpenStudy (victory):

use the inverse method to solve the following systems of equation x+y+z=14 2x-3y+4y=17 3x+y+2z=14

OpenStudy (anonymous):

first solve for one of your variables. lets solve for x first. so x=14-y-z now lets plug that into one of your other equations for x. so we get 2(14-y-z)-3y+4y=17 now lets combine our y terms. 28-2y-2z-3y+4y=17 simplifying we get y=11-2z